The sum of Mrs. Blacks age and her daughters age is 56 years. In 8 years, Mrs. Black will be twice as old as her daughter. How old is her daughter right now
Answer by tiffany222(56)

Let Mrs. Black's age be x
Let the daughter's age be y
x + y = 56
x + 8 = 2(y + 8)
x + 8 = 2y + 16
x - 2y = 8
Equations:
x + y = 56
x - 2y = 8
Subtract to eliminate x:
3y = 48
y = 16
Substitute y into one of the original equations to get x:
x + y = 56
x + 16 = 56
x = 40
Final Answer:
Mrs. Black's age = 40
Daughter's age = 16